OpenStudy (anonymous):

log base 3 of 9-log base 5 of 625 + log base 7 of 343

OpenStudy (gorv):

9=3^2 625=5^4 343=7^3

OpenStudy (gorv):

\[\log_{a} a=1\]

OpenStudy (gorv):

when base and number are same the log=1

OpenStudy (gorv):

\[\log_{3}3^2-\log_{5}5^4+\log_{7}7^3 \]

OpenStudy (gorv):

\[\log_{a} b^c=c *\log_{a}b \]

OpenStudy (anonymous):

thank you lot professor

OpenStudy (gorv):

can u do it ??

OpenStudy (anonymous):

sure sir thanks a lot

OpenStudy (anonymous):

i have another question will u please help me?

OpenStudy (anonymous):

yes sir i'm from india